OpenStudy (suwhitney):

You own an 5 year, 5 1/2% simple interest bearing note with a face value (principal) of $15,000. Three months before the due date of the note you need some money due to an unexpected emergency and decide to sell the note to a friend. If your friend wants to earn 8% simple interest on their investment, how much should your friend agree to pay you for the note? Round your answer off to the nearest whole dollar.

OpenStudy (anonymous):

holy crap thats a long problem

OpenStudy (suwhitney):

Haha, yes I know...

OpenStudy (anonymous):

@mayankdevnani a little help plz

OpenStudy (suwhitney):

I know that S = ? P = $15,000 R = 0.055 T = 5 years S = P + PRT So, S = 15,000 + (15,000)(0.055)(5) = 19125

OpenStudy (anonymous):

If FV is the final value of your interest bearing note after 3 months, then your friend wants to buy it at a price FP (friend price) such that \[FV = 1.08(FP)\]Solve for FP:\[FP = \frac{FV}{1.08}\]You are already given everything you need to caculate FV:\[FV=P(1+rt) = (15000)(1+(0.055)(5)) = 19125\]So \[FP = \frac{19125}{1.08} = 17708.33\]

OpenStudy (aum):

Shouldn't the 3 months figure into the calculation above?

OpenStudy (anonymous):

I don't see how it would. It would only affect your own profit/loss, not your friend's.

OpenStudy (suwhitney):

Thats what I was confused about because in a previous example she gave us the time that "friend 2" had the one was factored into the equation..

OpenStudy (aum):

The friend should buy it at such a price that in 3 months they would cash it for $19,125 yielding them an effective interest of 8%.

OpenStudy (aum):

x + x * 1/4 * .08 = 19,125 (3 months is 1/4th of a year). solve for x

OpenStudy (aum):

Friend's Principal = x simple interest = 8% = 0.08 time invested = 3 months = 1/4 year Final Amount = $19,125 Solve for x using the simple interest formula.

OpenStudy (suwhitney):

So, would it be $18,750?

OpenStudy (aum):

I think so.

OpenStudy (anonymous):

That would evaluate to x=18750. But if the friend spends 18750, then an 8% profit would bring it to 18750*1.08=20250, which is greater than the final value.

OpenStudy (aum):

No, but the friend will be cashing it in 3 months. What you calculated is after one year.

OpenStudy (aum):

The interest is 8% per annum. So you get the 8% only if you wait for 1 year.

OpenStudy (anonymous):

I guess it depends on if the 8% is supposed to be total profit or an annual rate. I interpreted it as a total profit, so 8% gained in just 3 months.

OpenStudy (aum):

8% simple interest implies 8% per year.

OpenStudy (suwhitney):

Thank you both for your help, I will let you know if I got it right :)